OpenStudy (anonymous):

George can rent a Cadillac from Gamma Car Rental for $29.17 per day plus 57 cents per mile. He can get the same car from Delta Car Rental for $50.23 per day plus 44 cents per mile. For what number of daily miles is Gamma cheaper?

OpenStudy (anonymous):

Can anyone help all I need is the answer cause I'm on a really tight deadline so please someone help.(p.s. you're helping me graduate)

OpenStudy (tpaulus):

Solve using a system of equations, g = 27.19 +57m d = 50.23 +44m find the point where they intersect...

OpenStudy (anonymous):

Let m be the number of miles driven. Solve the following equation for m:\[29+\frac{17}{100}+\frac{57}{100}m=50+\frac{23}{100}+\frac{44}{100}m\]m = 162 miles

OpenStudy (anonymous):

Ok thanks for explaining and more importantly the answer

OpenStudy (anonymous):

Your Welcome.

OpenStudy (anonymous):

By the way, I like to use fractions rather decimal numbers when possible. The result will always be an exact fraction or number.

OpenStudy (anonymous):

Aw ok I see